# ScribblePrompt: Fast and Flexible Interactive Segmentation for Any Biomedical Image

### v2

@@ Abstract
Biomedical image segmentation is a crucial part of both scientific research and clinical care. With enough labelled data, deep learning models can be trained to accurately automate specific biomedical image segmentation tasks. However, manually segmenting images to create training data is highly labor intensive and requires domain expertise. We present ScribblePrompt, a flexible neural network based interactive segmentation tool for biomedical imaging that enables human annotators to segment previously unseen structures using scribbles, clicks, and bounding boxes. Through rigorous quantitative experiments, we demonstrate that given comparable amounts of interaction, ScribblePrompt produces more accurate segmentations than previous methods on datasets unseen during training. In a user study with domain experts, ScribblePrompt reduced annotation time by 28% while improving Dice by 15% compared to the next best method. ScribblePrompt’s success rests on a set of careful design decisions. These include a training strategy that incorporates both a highly diverse set of images and tasks, novel algorithms for simulated user interactions and labels, and a network that enables fast inference. We showcase ScribblePrompt in an interactive demo and provide code athttps://scribbleprompt.csail.mit.edu
Figure 1: ScribblePrompt enables rapid iterative interactive segmentation of unseen tasks using bounding boxes, clicks, and scribbles. We show predictions from ScribblePrompt with iterative interaction steps on examples from datasets unseen during training. At each step, we visualize positive scribble and click inputs in green, negative scribble and click inputs in red, bounding box inputs in yellow, and the predicted segmentation in blue. Scribble thickness is enlarged for visual clarity. See Appendix for more examples.
@@ 1 Introduction
    is computationally efficient, enabling real-time inference, even on a single CPU.
To achieve this, we focused on design decisions aimed at ease of use and realistic interactions. To train our model, we started by gathering a large corpus of biomedical imaging datasets. We then designed task augmentation through synthesis strategies to encourage generalization to unseen tasks. Finally, we designed new interaction simulation strategies for training. The model itself uses an architecture optimized for efficient inference, to ensure real-time utility.
To evaluate ScribblePrompt, we compared its performance to that of existing interactive segmentation methods on datasets unseen during training. These experiments involved manually-collected scribbles, simulated interactions (Fig. 1), and a user study in which experienced annotators were asked to segment images. In the user study, ScribblePrompt reduced annotation time by 28% while increasing Dice by 15% compared to the next best method, the Segment Anything Model (SAM) [63]. Given similar amounts of interaction, ScribblePrompt achieved consistently better Dice scores than other methods.
We release an interactive tool, code, model weights, and our dataset of manually-collected scribbles at scribbleprompt.csail.mit.edu. We will also release new multi-annotator segmentation labels created using ScribblePrompt for select datasets.
@@ 2 Related Works
Foundation Models. Recent vision foundation models employ prompting to enable generalization to new tasks. These models, trained on large collections of (natural) image data, segment potentially-unseen structures specified by spatial prompts [63, 153], text descriptions [106, 63, 153], or examples [135, 136, 20, 27]. Some of these models can be used for interactive segmentation with spatial prompts.
Fine-tuning models initially developed using natural images is often unhelpful for specific tasks in the biomedical imaging domain, compared to training from scratch [107]. The Segment Anything Model (SAM) [63], a foundation model trained on natural images, performs well at biomedical image segmentation tasks with clear boundaries (e.g. organ segmentation in abdominal CT), but performs poorly on tasks and modalities involving more subtle delineations (e.g. deep structures in Brain MRI) [51, 92, 120, 45]. Several recent or concurrent papers fine-tune SAM for specific biomedical tasks or modalities [99, 75, 131, 60, 50, 145]. A few fine-tune SAM to segment biomedical images from multiple-modalities with bounding boxes [84], clicks [138] or both [24]. We show in our experiments that these models don’t perform well on many unseen tasks and are limiting in their required interactions.
@@
Scribbles are an intuitive form of interaction, however few interactive segmentation works employ them, because acquiring realistic scribbles for training and evaluation is challenging. Previous deep-learning models trained for scribble-based segmentation either relied on collecting large datasets of manually drawn scribbles [152, 73, 70] or simplistic approaches such as sampling random points [9, 6, 134] or Bezier curves [3]. One work has explored simulating more complex scribbles for natural image segmentation [23]. We build on these concepts to create a new scribble simulation engine for training and evaluation.
@@ 3 ScribblePrompt Approach
We present an interactive segmentation method that can generalize to new biomedical imaging modalities and regions of interest, while staying focused on practical usability. We describe the problem formulation, and present the important aspects of the ScribblePrompt framework: (i) simulation of realistic and flexible interactions during training, (ii) augmentation with synthetic labels during training to encourage generalization, and (iii) an efficient architecture for real-time interaction. We show an overview of training in Fig. 3 and 4.
@@ 3.1 Problem Formulation
We minimize the difference between the true segmentation $y^{t}$ and each of the iterative predictions $\hat{y}_{1},\dots,\hat{y}_{k}$,
@@ 3.2 Prompt Simulation
Scribbles. Given binary mask $y\in\{0,1\}^{h\times w}$, we simulate a scribble mask $s\in[0,1]^{h\times w}$ by first generating clean scribbles, and then corrupting them to account for user behavior and variability. We illustrate this process in Fig. 2 and Sec. B.1.1. To generate the clean scribbles, we start with one of these strategies:
@@
    Line Scribbles: We draw random lines by connecting two end points sampled from $\{(u,v)|y_{uv}=1\}$ (Fig. 8).
@@
    Centerline Scribbles: We simulate scribbles in the center of label $y$ using a thinning algorithm [147] that reduces the label to a 1-pixel wide skeleton (Fig. 9).
@@
    Contour Scribbles: We simulate a rough contour of the desired segmentation within the boundaries of the mask (Fig. 10). We first blur the mask to reduce the size of the label such that $\tilde{y}=\min(y,y\circ G_{k})$, where $G_{k}$ is a Gaussian blur kernel. Then we apply a threshold $\tilde{y}<h$ sampled in some intensity range $h\sim U[\tilde{y}_{min},\tilde{y}_{max}]$ and extract a contour inside the boundary of the mask.
@@
    Interior border region clicks: We sample clicks from a border region inside the boundary of the mask (Fig. 11). We first blur the mask $y$ to reduce the size of the label such that $\tilde{y}=\min(y,y\circ G_{k})$ where $G_{k}$ is a Gaussian blur kernel. We then sample click coordinates from $\{(u,v)|\tilde{y}_{uv}\in[a,b]\}$, where $a,b\sim U[\tilde{y}_{min},\tilde{y}_{max})$ are thresholds sampled in some intensity range.
@@
Iterative Training (Fig. 3). During the first step ($i=1$), we sample the combination of interactions and the number of initial positive and negative interactions $n_{pos},n_{neg}\sim U[n_{min},n_{max}]$. The initial interactions $u_{1}$ are simulated using the ground truth label $y^{t}$. In subsequent steps, correction scribbles or clicks are sampled from the error region $\varepsilon_{i-1}^{t}$ between the last prediction $\hat{y}^{t}_{i-1}$ and the ground truth $y^{t}$. Since a user can make multiple corrections in each step, we sample $n_{cor}\sim U[n_{min},n_{max}]$ corrections (scribbles or clicks) per step.
Figure 3: Training. We simulate $k$ consecutive steps of interactive segmentation. Given an image segmentation pair $(x^{t},y^{t})$, we first simulate a set of initial interactions $u_{1}$, which may contain bounding boxes, clicks, and/or scribbles. We predict segmentation $\hat{y}_{1}^{t}:=f_{\theta}(x^{t},u_{1},\hat{y}_{0}^{t})$ where the initial prediction $\hat{y}^{t}_{0}$ is set to zeros. In the second step, we simulate corrections using the error region $\varepsilon_{1}$ between the previous prediction $\hat{y}_{1}$ and ground truth $y$, and add them to the set of initial interactions $u_{1}$ to get $u_{2}$. We predict segmentation $\hat{y}_{2}:=f_{\theta}(x,u_{2},\hat{y}_{1})$ and repeat to produce a series of $k$ iterative predictions, $\hat{y}_{1},\dots,\hat{y}_{k}$. We learn $\theta$ to minimize $\sum_{i=1}^{k}\mathcal{L}_{seg}(y^{t},\hat{y}_{i})$, the sum of losses between the target segmentation $y^{t}$ and iterative predictions $\hat{y}_{1},\dots,\hat{y}_{k}$.
@@ 3.3 Data
We define a 2D segmentation task as a combination of dataset, axis (for 3D modalities), and label. For datasets with multiple segmentation labels, we consider each label separately as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice from each volume. We provide more details on the data in Appendix C.
@@
Synthetic Labels. To limit overfitting to specific segmentation tasks, we introduce a mechanism to generate synthetic labels (Fig. 4). During training, for a given sample $(x_{0},y_{0})$, with probability $p_{synth}$, we replace $y$ with a synthetic label $y_{synth}$. Given an image $x_{0}$, we generate a synthetic label $y_{synth}$ by applying a superpixel algorithm [33] with scale parameter $\lambda\sim U[1,\lambda_{max}]$ to partition the image into a multi-label mask of $k$ superpixels $z\in\{1,\dots,k\}^{n\times n}$, and then randomly select a superpixel $y_{synth}=\mathbbm{1}(z=c)$. We show examples in Fig. 14 and conduct experiments varying $p_{synth}$.
@@ 3.4 Network
Motivated by producing a practical tool, we primarily demonstrate ScribblePrompt using an efficient fully-convolutional architecture similar to a UNet [111]. We also demonstrate ScribblePrompt using a vision transformer-based architecture, SAM ViT-b [63]. We provide code and weights for both models at scribbleprompt.csail.mit.edu.
ScribblePrompt-UNet. We use an 8 layer CNN following a decoder-encoder structure similar to the popular UNet architecture [111] without Batch Norm. Each convolutional layer has 192 features and uses PReLu activation [44].
We encode each prompt type in an input channel. The input to ScribblePrompt-UNet has size $5\times h\times w$ consisting of the input image $x^{t}$, bounding box encoding, positive click/scribble encoding, negative click/scribble encoding, and the logits of the previous prediction $\hat{y}^{t}_{i-1}$. We encode bounding boxes in a binary mask that is 1 inside the box(es) and 0 everywhere else. We encode positive and negative clicks using binary masks, and combine them with the positive and negative scribble masks.
ScribblePrompt-SAM. We take the smallest SAM model (ViT-b) and fine-tune the decoder. SAM takes bounding box and click inputs as lists of ${(x,y)}$ coordinates, and the low-resolution logits of the previous prediction $\hat{y}_{i-1}$. To adapt SAM for scribbles, we consider each positive/negative scribbled pixel as a positive/negative click.
Loss. We minimize eq. (1) where $\mathcal{L}_{Seg}$ is a linear combination of soft Dice Loss [32] and Focal Loss [74]. In preliminary experiments (Sec. B.2), training with Soft Dice Loss [32] alone or a linear combination of Soft Dice Loss and Binary-Cross Entropy Loss resulted in slightly lower Dice scores.
@@ 4 Experimental Setup
MedSAM [84]: MedSAM is a SAM ViT-B model fine-tuned with bounding box prompts on a collection of biomedical imaging datasets containing 1.5M image segmentation pairs. We evaluate MedSAM with bounding boxes only, because we found it to perform poorly when given point or mask prompts (Fig. 24).
MIDeepSeg [82]: MIDeepSeg is an interactive segmentation framework designed to generalize to unseen tasks on medical images. MIDeepSeg takes interior margin points (positive clicks) as initial inputs. It crops the image based on those points and uses a CNN to make an initial prediction. Given additional clicks, the prediction can be refined using conditional random fields. We evaluate the pre-trained model, which was developed on placenta T2 MRI.
@@ 5 Evaluation
Figure 5: Example predictions. SP = ScribblePrompt. Top: predictions after one step of manual scribbles. Bottom: predictions after five steps of simulated interactions (one center click followed by one center correction click per step). See the Appendix E,F for more examples.
@@ 5.1 Manual Scribbles
Setup. We evaluate on two datasets of manual scribbles. First, we collected a dataset (MedScribble) of manual scribbles from three annotators for seven segmentation tasks from unseen datasets, with a total of 31 examples. The annotators were shown five training examples with the ground truth segmentation per task and instructed to draw positive and negative scribbles to indicate the region of interest on new images. We provide more details on MedScribble in Appendix E. Second, we used 380 slices from the ACDC dataset [14], which has scribble annotations for three labels and background.
@@
Results. ScribblePrompt-UNet and ScribblePrompt-SAM produce the most accurate segmentations in a single step of manual scribbles on both our manual scribbles dataset and the ACDC scribbles dataset (Tab. 1). We show example predictions in Fig. 5 and Appendix E.
@@
Table 1: Manual scribbles. We report mean Dice and HD95 with 95% CI of segmentations predicted from manually-collected scribbles. Best and second best are highlighted.
MedScribble (n=93) ACDC (n=2,130)

### v3

@@ Abstract
Biomedical image segmentation is a crucial part of both scientific research and clinical care. With enough labelled data, deep learning models can be trained to accurately automate specific biomedical image segmentation tasks. However, manually segmenting images to create training data is highly labor intensive and requires domain expertise. We present ScribblePrompt, a flexible neural network based interactive segmentation tool for biomedical imaging that enables human annotators to segment previously unseen structures using scribbles, clicks, and bounding boxes. Through rigorous quantitative experiments, we demonstrate that given comparable amounts of interaction, ScribblePrompt produces more accurate segmentations than previous methods on datasets unseen during training. In a user study with domain experts, ScribblePrompt reduced annotation time by 28% while improving Dice by 15% compared to the next best method. ScribblePrompt’s success rests on a set of careful design decisions. These include a training strategy that incorporates both a highly diverse set of images and tasks, novel algorithms for simulated user interactions and labels, and a network that enables fast inference. We showcase ScribblePrompt in an interactive demo, provide code, and release a dataset of scribble annotations at https://scribbleprompt.csail.mit.edu
Keywords:
Interactive Segmentation Biomedical Imaging Scribbles
@@ 1 Introduction
Figure 1: ScribblePrompt enables rapid iterative interactive segmentation of unseen tasks using bounding boxes, clicks, and scribbles. We show predictions from ScribblePrompt with iterative interaction steps on examples from datasets unseen during training. At each step, we visualize positive scribble and click inputs in green, negative scribble and click inputs in red, bounding box inputs in yellow, and the predicted segmentation in blue. Scribble thickness is enlarged for visual clarity. See Supplementary Material for more examples.
@@
    is computationally efficient, enabling fast inference, even on a single CPU.
To achieve this, we focused on design decisions aimed at ease of use and realistic interactions. To train our model, we started by gathering a large corpus of biomedical imaging datasets. We then designed task augmentation through synthesis strategies to encourage generalization to unseen tasks. Finally, we designed new interaction simulation strategies for training. The model itself uses an architecture optimized for efficient inference.
To evaluate ScribblePrompt, we compared its performance to that of existing interactive segmentation methods on datasets unseen during training. These experiments involved manually-collected scribbles, simulated interactions (Fig. 1), and a user study in which experienced annotators were asked to segment images. In the user study, ScribblePrompt reduced annotation time by 28% while increasing Dice by 15% compared to the next best method, the Segment Anything Model (SAM) [64]. Given similar amounts of interaction, ScribblePrompt achieved consistently better Dice scores than other methods. We release an interactive tool, code, model weights, and our dataset of manually-collected scribbles at https://scribbleprompt.csail.mit.edu.
@@ 2 Related Works
Foundation Models. Recent vision foundation models employ prompting to enable generalization to new tasks. These models, trained on large collections of (natural) image data, segment potentially-unseen structures specified by spatial prompts [64, 159], text [112, 64, 159], or examples [140, 141, 20, 27]. Some of these models can be used for interactive segmentation with spatial prompts.
Fine-tuning models initially developed using natural images is often unhelpful for specific tasks in the biomedical imaging domain, compared to training from scratch [113]. The Segment Anything Model (SAM) [64], a foundation model trained on natural images, performs well at biomedical image segmentation tasks with clear boundaries (e.g. organs in abdominal CT), but performs poorly on tasks and modalities involving more subtle delineations (e.g. deep structures in Brain MRI) [53, 98, 126, 47]. Several recent or concurrent papers fine-tune SAM for specific biomedical tasks or modalities [105, 78, 136, 61, 52, 150]. A few fine-tune SAM to segment medical images from multiple-modalities with bounding boxes [89], clicks [143] or both [24]. We show in our experiments that these models don’t perform well on many unseen tasks and are limiting in their required interactions.
@@
Scribbles are an intuitive form of interaction, however few interactive segmentation works employ them, because acquiring realistic scribbles for training and evaluation is challenging. Previous deep-learning models trained for interactive scribble-based segmentation either relied on collecting large datasets of manually drawn scribbles [158, 76, 71] or simplistic approaches such as sampling random points [9, 6, 138] or Bezier curves [3]. One work has explored simulating more complex scribbles for interactive segmentation of natural images [23]. We build on these concepts to create a new scribble simulation engine for training and evaluation.
Scribble-supervised learning methods use scribble annotations as supervision to train automatic segmentation models for predicting segmentation given only an input image [76, 74, 151, 86, 40]. However, these methods require the manual scribble-annotation of many training images and retraining for each new task. In contrast, ScribblePrompt is trained on a large corpus of datasets once and then can be used to perform new segmentation tasks at inference time without retraining, using scribbles as input.
@@ 3 ScribblePrompt Approach
We present an interactive segmentation method that can generalize to new biomedical imaging modalities and regions of interest, while staying focused on practical usability. We describe the problem formulation, and present the important aspects of the ScribblePrompt framework: (i) simulation of realistic interactions during training, (ii) augmentation with synthetic labels during training to encourage generalization, and (iii) an efficient architecture for fast inference. We show an overview of training in Fig. 2 and 4.
@@ 3.1 Problem Formulation
We minimize the difference between the true segmentation $y^{t}$ and each of the $k$ iterative predictions $\hat{y}_{1},\dots,\hat{y}_{k}$,
@@
Figure 2: Training. We simulate $k$ consecutive steps of interactive segmentation. Given an image segmentation pair $(x^{t},y^{t})$, we first simulate a set of initial interactions $u_{1}$, which may contain bounding boxes, clicks, and/or scribbles. We predict segmentation $\hat{y}_{1}^{t}:=f_{\theta}(x^{t},u_{1},\hat{y}_{0}^{t})$ where the initial prediction $\hat{y}^{t}_{0}$ is set to zeros. In the second step, we simulate corrections using the error region $\varepsilon_{1}^{t}$ between the previous prediction $\hat{y}_{1}^{t}$ and ground truth $y^{t}$, and add them to the set of initial interactions $u_{1}$ to get $u_{2}$. We predict segmentation $\hat{y}_{2}^{t}:=f_{\theta}(x^{t},u_{2},\hat{y}_{1}^{t})$ and repeat to produce a series of predictions, $\hat{y}_{1}^{t},\dots,\hat{y}_{k}^{t}$. We learn $\theta$ to minimize $\sum_{i=1}^{k}\mathcal{L}_{seg}(y^{t},\hat{y}_{i}^{t})$, the sum of losses between the target segmentation $y^{t}$ and iterative predictions $\hat{y}_{1}^{t},\dots,\hat{y}_{k}^{t}$.
@@ 3.2 Prompt Simulation
Scribbles. Given binary mask $y\in\{0,1\}^{h\times w}$, we simulate a scribble mask $s\in[0,1]^{h\times w}$ by first generating clean scribbles, and then corrupting them to account for user behavior and variability. We illustrate this process in Supplementary Material Sec. 0.B.1.1. To generate the clean scribbles, we start with one of these strategies:
@@
    Line Scribbles: We draw random lines by connecting two end points sampled from $\{(u,v)|y_{uv}=1\}$.
@@
    Centerline Scribbles: We simulate scribbles in the center of label $y$ using a thinning algorithm [153] that reduces the label to a 1-pixel wide skeleton.
@@
    Contour Scribbles: We simulate a rough contour of the desired segmentation within the boundaries of the mask. We first blur the mask to reduce the size of the label such that $\tilde{y}=\min(y,y\circ G_{k})$, where $G_{k}$ is a Gaussian blur kernel. Then we apply a threshold $\tilde{y}<h$ sampled in some intensity range $h\sim U[\tilde{y}_{min},\tilde{y}_{max}]$ and extract a contour inside the boundary of the mask.
@@
    Interior border region clicks: We sample clicks from a border region inside the boundary of the mask. We first blur the mask $y$ to reduce the size of the label such that $\tilde{y}=\min(y,y\circ G_{k})$ where $G_{k}$ is a Gaussian blur kernel. We then sample click coordinates from $\{(u,v)|\tilde{y}_{uv}\in[a,b]\}$, where $a,b\sim U[\tilde{y}_{min},\tilde{y}_{max})$ are thresholds sampled in some intensity range.
@@
Iterative Training. During the first step ($i=1$), we sample the combination of interactions and the number of initial positive and negative interactions $n_{pos},n_{neg}\sim U[n_{min},n_{max}]$. The initial interactions $u_{1}$ are simulated using the ground truth label $y^{t}$. In subsequent steps, correction scribbles or clicks are sampled from the error region $\varepsilon_{i-1}^{t}$ between the last prediction $\hat{y}^{t}_{i-1}$ and the ground truth $y^{t}$. Since a user can make multiple corrections in each step, we sample $n_{cor}\sim U[n_{min},n_{max}]$ corrections (scribbles or clicks) per step.
@@ 3.3 Data
We define a 2D segmentation task as a combination of dataset, axis (for 3D modalities), and label. For datasets with multiple segmentation labels, we consider each label separately as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice from each volume. We provide more details on the data in Supplementary Material Sec. 0.C.
@@
Synthetic Labels. To limit overfitting to specific segmentation tasks, we introduce a mechanism to generate synthetic labels (Fig. 4). During training, for a given sample $(x_{0},y_{0})$, with probability $p_{synth}$, we replace $y$ with a synthetic label $y_{synth}$. Given an image $x_{0}$, we generate a synthetic label $y_{synth}$ by applying a superpixel algorithm [33] with scale parameter $\lambda\sim U[1,\lambda_{max}]$ to partition the image into a multi-label mask of $k$ superpixels $z\in\{1,\dots,k\}^{n\times n}$, and then randomly select a superpixel $y_{synth}=\mathbbm{1}(z=c)$. We conduct experiments varying $p_{synth}$ and show examples in Supplementary Material Sec. 0.B.3.
@@ 3.4 Network
Motivated by producing a practical tool, we primarily demonstrate ScribblePrompt using an efficient fully-convolutional architecture similar to a UNet [117]. We also demonstrate ScribblePrompt using a vision transformer architecture [64].
ScribblePrompt-UNet: We use an 8 layer CNN following a decoder-encoder structure similar to the popular UNet architecture [117] without Batch Norm. Each convolutional layer has 192 features and uses PReLu activation [46].
ScribblePrompt-SAM: We take the smallest SAM model (ViT-b)  [64] and fine-tune the decoder. SAM takes bounding box and click inputs as lists of ${(x,y)}$ coordinates, and the low-resolution logits of the previous prediction $\hat{y}_{i-1}$. To adapt SAM for scribbles, we consider each scribbled pixel as a click.
Loss. We minimize eq. (1) where $\mathcal{L}_{Seg}$ is a linear combination of soft Dice Loss [32] and Focal Loss [77]. In preliminary experiments (Supplementary Material Sec. 0.B.2), training with Soft Dice Loss alone or a linear combination of Soft Dice Loss and Binary-Cross Entropy Loss resulted in slightly lower Dice scores.
@@ 4 Experimental Setup
MedSAM [89]: MedSAM is a SAM ViT-B model fine-tuned with bounding box prompts on a collection of biomedical imaging datasets containing 1.5M image segmentation pairs. We evaluate MedSAM with bounding boxes only, because we found it to perform poorly when given point or mask prompts.
MIDeepSeg [88]: MIDeepSeg is an interactive segmentation framework designed to generalize to unseen tasks on medical images. MIDeepSeg takes interior margin points (positive clicks) as initial inputs, crops the image based on those points, and uses a CNN to make an initial prediction. Given additional clicks, the prediction can be refined using conditional random fields. We evaluate the pre-trained model, which was developed on placenta T2 MRI.
@@ 5.1 Manual Scribbles
Setup. We evaluate on two datasets of manual scribbles. First, we collected a dataset (MedScribble) of manual scribbles from three annotators for seven segmentation tasks from unseen datasets, with a total of 31 examples. The annotators were shown five training examples with the ground truth segmentation per task and instructed to draw positive and negative scribbles to indicate the region of interest on new images. We provide more details on MedScribble and release the data in Supplementary Material Sec. 0.E.1. Second, we used 380 slices from the ACDC dataset [14], which has scribble annotations for three labels and background.
@@
Results. ScribblePrompt-UNet and ScribblePrompt-SAM produce the most accurate segmentations in a single step of manual scribbles on both our manual scribbles dataset and the ACDC scribbles dataset (Tab. 1). We show example predictions in Fig. 5 and additional comparisons to scribble-supervised learning in Supplementary Material Sec. 0.E.3.
@@
Table 1: Manual scribbles. Mean Dice and HD95 with 95% CI of segmentations predicted from manually-collected scribbles. Best and second best are highlighted.
MedScribble (n=93) ACDC (n=1,140)
@@
Figure 5: Example predictions. SP = ScribblePrompt. Top: predictions after one step of manual scribbles. Bottom: predictions after five steps of simulated interactions (one center click followed by one center correction click per step).